Can someone help me with this problem?
How do you write 1/8 as a percent.
Thanks,
Sher

Answer by tutorcecilia(2144)

1/8 can be thought of as 1 divided by 8. Think of the fraction bar as a division symbol in which you divide the numerator by the denominator.
1/8 = 0.125
.
Example: 8/1 = 8 divided by 1 = 8
Example: 3/36 = 3 divided by 36 = .08333
Example: 36/3 = 36 divided by 3 = 12
.
Now to change any of these decimals into a percentage, just multiply by 100%.
1/8 = 0.125 X 100% = 12.5%
3/36 = .0833 x 100% = 8.33%
36/3 = 12 x 100% =1200%